# Wild Berry image dataset collected in Finnish forests and peatlands using drones

### v1

@@ Abstract
Berry picking has long-standing traditions in Finland, yet it is challenging and can potentially be dangerous. The integration of drones equipped with advanced imaging techniques represents a transformative leap forward, optimising harvests and promising sustainable practices. We propose WildBe, the first image dataset of wild berries captured in peatlands and under the canopy of Finnish forests using drones. Unlike previous and related datasets, WildBe includes new varieties of berries, such as bilberries, cloudberries, lingonberries, and crowberries, captured under severe light variations and in cluttered environments. WildBe features 3,516 images, including a total of 18,468 annotated bounding boxes. We carry out a comprehensive analysis of WildBe using six popular object detectors, assessing their effectiveness in berry detection across different forest regions and camera types. We will release WildBe publicly. ††This work was supported by the EU Horizon Europe project FEROX under Grant 101070440.
1 Introduction
@@ 2 Berries
Lingonberry (Vaccinium vitis-idaea) and bilberry (Vaccinium myrtillus) have adapted to a wide range of different site and land types in coniferous ecosystems and, as a result, are widely distributed across Europe and northern Asia. In Finland, bilberry is typical and abundant, especially in conifer heath forests of medium site fertility (i.e. mesic heath forests). Lingonberry is most typical in light pine-dominated dryish (sub-xeric) heath forests. Both species also occur and produce yields in marginal types of forest (e.g. fell forests), and on pristine and drained peatland sites [16]. The leaves of the lingonberry plant are small, oval-shaped, and have smooth edges. They are dark green, glossy on the top, and can sometimes exhibit a slight reddish tint along the edges. The lingonberry plant retains its leaves, which even survive through the winter. Lingonberry flowers are bell-shaped, white to pale pink, and grow in small clusters. They are small and round, with a diameter of about 5-8 mm. Berries are initially light green, turning red upon ripening. Lingonberry plants are low-growing shrubs, typically reaching a height of 10-30 cm. The plants have a woody stem and can spread over a wide area. Differently, the leaves of the bilberry plant are small, oval to elliptical, and have finely serrated edges. They are bright green and soft, growing along the slender, green branches. The flowers are small, usually pink in colour, and their shape is ball-like. The berries are round and small, about 5-8 mm in diameter. They are either dark blue and waxy or black and shiny. Bilberry plants are low-growing shrubs, usually about 20-40 cm in height. They form dense, twiggy clumps and can spread over the ground in extensive patches, particularly in undisturbed habitats.
@@ 3.2 Cameras
WildBe features different sensors, ranging from mobile, action, to drone cameras. Data includes images captured with the Xiaomi 12X smartphone, equipped with a 50MP Sony IMX766 camera [8]. The DJI Mini 2 drone offers high-quality imagery with its 12MP camera, adaptable to various environmental conditions [15]. The Mavic 2 Pro mounts the Hasselblad L1D-20c camera with a 1-inch CMOS sensor, known to excel in low-light environments, thanks to its extended ISO range and enhanced dynamic range [4]. The DJI Mavic 3M mounts a 20MP RGB multispectral camera, already employed in agriculture applications [14]. The GoPro 11, mounted on our custom drone named Scout v2, provides action-oriented imaging with its 27-megapixel sensor and HyperSmooth 5.0 video stabilisation technology. To the best of our knowledge, the GoPro 11 has not yet been utilised in agriculture applications; however, prior versions of the camera have demonstrated performance comparable to high-end models [3]. Table 1 summaries the cameras used.
@@ 4 Dataset
WildBe was collected in the forests of Ilomantsi (Finland) in July 2023. It comprises 3,516 images, extracted from 555 raw images selected from different drone models and devices (Sec. 4.2), and encompasses variations of cameras, albedos, forests, terrains, heights, perspective changes, and includes a total of 18,336 berries manually annotated as bounding boxes. Berry species include bilberries, cloudberries, crowberries, and lingonberries, the last at different ripening stages. To provide a detailed reference of where the data was captured, WildBe incorporates metadata for each image, sourced from various devices (Tab. 1). This metadata includes: timestamp of each captured image, geographical coordinates, altitude, and camera details. While DJI drones automatically capture this metadata, images from Xiaomi and GoPro cameras include only the capture date.
@@ 4.2 Pre-processing and annotations
Given the original images captured by drones, we create smaller versions through cropping to enhance the efficiency of object detection algorithms. The cameras we use have different resolutions (Tab. 1). We select 555 original images from the collection, taking into account factors such as species, device, and lighting conditions, and post-process them into non-overlapping image crops of size 528396 pixels. In instances where image crops111Hereafter, we refer to these as ‘images’ for simplicity. are smaller, we rescale them using bilinear interpolation. These image crops are manually annotated with bounding boxes in the YOLO format (i.e., (). The classes correspond to cloudberries, lingonberries, crowberries, and bilberries. For each image, bounding boxes are drawn as tightly as possible around the berries to minimise the inclusion of background elements. Since bilberries and crowberries are both dark in colour and difficult to distinguish, annotators rely on the shape of the leaves to differentiate them (bilberry leaves are wide-open, whereas crowberry leaves are needle-shaped). WildBe also accounts for varying ripeness levels of lingonberries, from green to reddish hues, treating all as equivalent representations of the same species. Although less represented in numbers, cloudberries can be easily identified based by their distinctive yellow colour.
@@ 4.3 Annotation statistics
Fig. 2 illustrates the proportion of annotations for each class within the dataset (on the left-hand side) and the distribution of bounding boxes across the images (on the right-hand side). We can observe that the majority of the annotations are for bilberries, while cloudberries constitute the minority. However, due to the distinctive yellow appearance of cloudberries, experiments demonstrate that detectors are particularly effective in identifying them. Moreover, it is also evident that the annotations are evenly distributed throughout the dataset, which aids in preventing the development of biases during the training of algorithms.
[…]
@@ 5.2 Detectors
We compare six popular object detectors: Faster R-CNN (2015) [13], VarifocalNet (2021) [21], GLIP (2022) [9], DINO (2023) [20], ObjectBox (2022) [19], and YOLOv8 (2023) [7]. We use the MMDetection open source library [5] for the implementation of the first four methods, whereas we employ the authors’ code for ObjectBox and YOLOv8. For a fair comparison, we select the algorithm’s backbones to have similar number of parameters.
5.3 Quantitative results
@@
Cross-camera transfer learning. We evaluate the ability of algorithms to transfer across different cameras. Tab. 4 displays the detection results when algorithms are trained with data captured by four cameras and then tested on a different camera (Tab. 1). In the multi-class scenario, GLIP outperforms other algorithms in terms of Avg, followed by DINO and VerifocalNet. Faster-RCNN reports the lowest average in most cases. In the single-class scenario, GLIP again outperforms in transfer learning cases, except in the case of the DJI Mini 2, where DINO performs best. When the number of test instances is imbalanced (e.g., in the case of DJI Mavic 2 Pro), APs vary significantly, yet the average APs does not accurately reflect each algorithm’s performance. In such cases, WAvg is a more rational metric to prevent class bias. Comparing WAvg amongst the four classes with AP of a single class, we observe three behaviours. Firstly, in the case of a single class, AP remains higher than the multiclass WAvg (e.g., Xiaomi 12X), as the single-class AP reflects only a localisation task, while WAvg scores also include an additional classification task that may hinder performance. Secondly, it is often observed that the single-class APs and the multiclass WAvg are comparable (e.g., DJI Mavic 2 Pro), indicating that the algorithms perform well both in classification and localisation. Thirdly, a higher WAvg than single-class AP suggests that classification can enhance the algorithm’s learning in localisation abilities (e.g., GoPro 11). Overall, considering both the single class APs as well as the multi-class WAvg, we observe that DJI Mavic 2 Pro is the easiest sensor to transfer to, probably since its test set is dominated by samples from the Bilberry class that report the highest score amongst the four classes. On the other hand, DJI Mavic 3M is the hardest sensor to transfer to as it contains several Crowberry and Lingonberry instances that are the most difficult to handle.
@@
Ground-truth
Cross-dataset transfer learning. We evaluate the algorithms’ ability to generalise across datasets, i.e., trained on WildBe and tested on CRAID [2]. CRAID contains only cranberries, which are not included in WildBe. Moreover, CRAID is captured with a different sensor, in various geographical locations, and under different acquisition conditions. Tab. 5 presents the results. Consistent with previous results, GLIP performs the best. In contrast, DINO, which ranks second-best in prior evaluations, exhibits a significant performance drop on CRAID. Overall, this cross-dataset experiment demonstrates a notable decline in performance for all the algorithms, attributable to the severe domain discrepancy between the training set (WildBe) and the test set.
@@ 6 Conclusions
We presented WildBe, a new image dataset of wild berries collected in Finnish forests using drone technology, marking a significant advancement for the automation of berry picking and agricultural practices. Unlike existing datasets, such as CRAID [2], which focuses on cranberry cultivation fields, WildBe encompasses a diverse range of wild berries (bilberries, cloudberries, crowberries, and lingonberries) captured in challenging conditions of forest undercanopies. WildBe features 3,516 images with 18,468 annotated bounding boxes, WildBe provides a rich resource for developing and testing advanced object detection algorithms. We comprehensively analysed six popular object detectors (Faster R-CNN, VarifocalNet, GLIP, DINO, ObjectBox, and YOLOv8) to assess the dataset’s utility in enhancing the performance and generalisation ability of detection models across varied forest regions and camera types. One limitation of our dataset is the annotations for the bilberry species; that is, we annotated bilberries and bog bilberries as the same class (species). These two species are very similar to each other, and only experts can accurately distinguish between them. As future work, we plan to involve experts for such fine-grained annotation to add more value to WildBe.

### v3

@@ Abstract
Berry picking has long-standing traditions in Finland, yet it is challenging and can potentially be dangerous. The integration of drones equipped with advanced imaging techniques represents a transformative leap forward, optimising harvests and promising sustainable practices. We propose WildBe, the first image dataset of wild berries captured in peatlands and under the canopy of Finnish forests using drones. Unlike previous and related datasets, WildBe includes new varieties of berries, such as bilberries, cloudberries, lingonberries, and crowberries, captured under severe light variations and in cluttered environments. WildBe features 3,516 images, including a total of 18,336 annotated bounding boxes. We carry out a comprehensive analysis of WildBe using six popular object detectors, assessing their effectiveness in berry detection across different forest regions and camera types. WildBe is publicly available on HuggingFace (https://huggingface.co/datasets/FBK-TeV/WildBe).
Keywords:
Object detection Dataset Drone imagery Agritech1 Introduction
@@ 2 Berries
Lingonberry (Vaccinium vitis-idaea) and bilberry (Vaccinium myrtillus) have adapted to a wide range of different site and land types in coniferous ecosystems and, as a result, are widely distributed across Europe and northern Asia. In Finland, bilberry is typical and abundant, especially in conifer heath forests of medium site fertility (e.g., mesic heath forests). Lingonberry is most typical in light pine-dominated dryish (sub-xeric) heath forests. Both species also occur and produce yields in marginal types of forest (e.g., fell forests), and on pristine and drained peatland sites [16]. The leaves of the lingonberry plant are small, oval-shaped, and have smooth edges. They are dark green, glossy on the top, and can sometimes exhibit a slight reddish tint along the edges. The lingonberry plant retains its leaves, which even survive through the winter. Lingonberry flowers are bell-shaped, white to pale pink, and grow in small clusters. They are small and round, with a diameter of about 5-8 mm. Berries are initially light green, turning red upon ripening. Lingonberry plants are low-growing shrubs, typically reaching a height of 10-30 cm. The plants have a woody stem and can spread over a wide area. Differently, the leaves of the bilberry plant are small, oval to elliptical, and have finely serrated edges. They are bright green and soft, growing along the slender, green branches. The flowers are small, usually pink in colour, and their shape is ball-like. The berries are round and small, about 5-8 mm in diameter. They are either dark blue and waxy or black and shiny. Bilberry plants are low-growing shrubs, usually about 20-40 cm in height. They form dense, twiggy clumps and can spread over the ground in extensive patches, particularly in undisturbed habitats.
@@ 3.2 Cameras
WildBe features different sensors, ranging from mobile, action, to drone cameras. Data also includes images captured with the Xiaomi 12X smartphone, equipped with a 50MP Sony IMX766 camera [7]. The DJI Mini 2 drone offers high-quality imagery with its 12MP camera, adaptable to various environmental conditions [15]. The Mavic 2 Pro mounts the Hasselblad L1D-20c camera with a 1-inch CMOS sensor, known to excel in low-light environments, thanks to its extended ISO range and enhanced dynamic range [3]. The DJI Mavic 3M mounts a 20MP RGB multispectral camera, already employed in agriculture applications [14]. The GoPro 11, mounted on our custom drone named Scout v2, provides action-oriented imaging with its 27-megapixel sensor and HyperSmooth 5.0 video stabilisation technology. To the best of our knowledge, the GoPro 11 has not yet been utilised in agriculture applications; however, prior versions of the camera have demonstrated performance comparable to high-end models [2]. Table 1 summaries the cameras used.
@@ 4 Dataset
WildBe was collected in the forests of Ilomantsi (Finland) in July 2023. It comprises 3,516 images, extracted from 555 raw images selected from different drone models and devices (Sec. 4.2), and encompasses variations of cameras, albedos, forests, terrains, heights, perspective changes, and includes a total of 18,336 berries manually annotated as bounding boxes. Berry species include bilberries, cloudberries, crowberries, and lingonberries, the last at different ripening stages. To provide a detailed reference of where the data was captured, WildBe includes metadata for each image, sourced from various devices (Tab. 1). This metadata includes: timestamp of each captured image, geographical coordinates, altitude, and camera details. While DJI drones automatically capture this metadata, images from Xiaomi and GoPro cameras include only the capture date.
@@ 4.2 Pre-processing and annotations
Given the original images captured by drones, we create smaller versions through cropping to enhance the efficiency of object detection algorithms. The cameras we use have different resolutions (Tab. 1). We select 555 original images from the collection, taking into account factors such as species, device, and lighting conditions, and post-process them into non-overlapping image crops of size 528396 pixels. In instances where image crops111Hereafter, we refer to these as ‘images’ for simplicity. are smaller, we rescale them using bilinear interpolation. These image crops are manually annotated with bounding boxes in the YOLO format (i.e., (). The classes correspond to cloudberries, lingonberries, crowberries, and bilberries. For each image, bounding boxes are drawn as tightly as possible around the berries to minimise the inclusion of background. Since bilberries and crowberries are both dark in colour and difficult to distinguish, annotators rely on the shape of the leaves to differentiate them (bilberry leaves are wide-open, whereas crowberry leaves are needle-shaped). WildBe also accounts for varying ripeness levels of lingonberries, from green to reddish hues, treating all as equivalent representations of the same species. Although less represented in numbers, cloudberries can be easily identified based by their distinctive yellow colour.
[…]
@@ 4.3 Annotation statistics
Fig. 2 illustrates the proportion of annotations for each class within the dataset (on the left-hand side) and the distribution of bounding boxes across the images (on the right-hand side). We can observe that the majority of the annotations are for bilberries, while cloudberries constitute the minority. However, due to the distinctive yellow appearance of cloudberries, experiments demonstrate that detectors are particularly effective in identifying them. Moreover, it is also evident that the annotations are evenly distributed across the images, which aids in preventing the development of biases during the training of algorithms.
@@ 5.2 Detectors
We compare six popular object detectors: Faster R-CNN (2015) [13], VarifocalNet (2021) [21], GLIP (2022) [8], DINO (2023) [20], ObjectBox (2022) [19], and YOLOv8 (2023) [6]. We use the MMDetection open source library [4] for the implementation of the first four methods, whereas we employ the authors’ code for ObjectBox and YOLOv8. For a fair comparison, we select the algorithm’s backbones to have similar number of parameters: we use ResNet50 for Faster R-CNN, VarifocalNet, and DINO; DarkNet for ObjectBox, and YOLOv8, and Swin-T for GLIP.
@@ 5.3 Quantitative results
Cross-camera transfer learning. We evaluate the ability of algorithms to transfer across different cameras. Tab. 4 displays the detection results when algorithms are trained with data captured by four cameras and then tested on a different camera (Tab. 1). In the multi-class scenario, GLIP outperforms other algorithms in terms of Avg, followed by DINO and VerifocalNet. Faster-RCNN reports the lowest average in most cases. In the single-class scenario, GLIP again outperforms in transfer learning cases, except in the case of the DJI Mini 2, where DINO performs best. When the number of test instances is imbalanced (e.g., in the case of DJI Mavic 2 Pro), APs vary significantly, yet the average APs does not accurately reflect each algorithm’s performance. In such cases, WAvg is a more rational metric to prevent class bias. Comparing WAvg amongst the four classes with AP of a single class, we observe three behaviours. Firstly, in the case of a single class, AP remains higher than the multiclass WAvg (e.g., Xiaomi 12X), as the single-class AP reflects only a localisation task, while WAvg scores also include an additional classification task that may hinder performance. Secondly, it is often observed that the single-class APs and the multiclass WAvg are comparable (e.g., DJI Mavic 2 Pro), indicating that the algorithms perform well both in classification and localisation. Thirdly, a higher WAvg than single-class AP suggests that classification can enhance the algorithm’s learning in localisation abilities (e.g., GoPro 11). Overall, considering both the single class APs as well as the multi-class WAvg, we observe that DJI Mavic 2 Pro is the easiest sensor to transfer to, probably since its test set is dominated by samples from the Bilberry class that report the highest score amongst the four classes. Oppositely, DJI Mavic 3M is the hardest sensor to transfer to as it contains several Crowberry and Lingonberry instances that are the most difficult to handle.
@@
Ground truth
Cross-dataset transfer learning. We evaluate the algorithms’ ability to generalise across datasets, i.e., trained on WildBe and tested on the 702 images of the test set of CRAID [1]. CRAID contains only cranberries, which are not included in WildBe. Moreover, CRAID is captured with a different sensor, in various geographical locations, and under different acquisition conditions. Tab. 5 presents the results. Consistent with previous results, GLIP performs the best. In contrast, DINO, which ranks second-best in prior evaluations, exhibits a significant performance drop on CRAID. Overall, this cross-dataset experiment shows a notable decline in performance for all the algorithms, attributable to the severe domain discrepancy between the training set (WildBe) and the test set.
@@ 6 Conclusions
We presented WildBe, a new image dataset of wild berries collected in Finnish forests using drone technology, marking a significant advancement for the automation of berry picking and agricultural practices. Unlike existing datasets, such as CRAID [1], which focuses on cranberry cultivation fields, WildBe encompasses a diverse range of wild berries (bilberries, cloudberries, crowberries, and lingonberries) captured in challenging conditions of forest undercanopies. WildBe features 3,516 images with 18,336 annotated bounding boxes, WildBe provides a rich resource for developing and testing advanced object detection algorithms. We comprehensively analysed six popular object detectors (Faster R-CNN, VarifocalNet, GLIP, DINO, ObjectBox, and YOLOv8) to assess the dataset’s utility in enhancing the performance and generalisation ability of detection models across varied forest regions and camera types. One limitation of our dataset is the annotations for the bilberry species; i.e., we annotated bilberries and bog bilberries as the same class (species). These two species are very similar to each other, and only experts can accurately distinguish between them. As future work, one can explore domain adaptation techniques to improve cross-dataset transfer learning [10]. Moreover, we plan to involve experts for such fine-grained annotation to add more value to WildBe.
Acknowledgement. This work was supported by the EU Horizon Europe project FEROX under Grant n. 101070440.